# Geometry is Wavy: Curvature Wave Equations for Generic Affine Connections

### v1

@@ A.2 Extraction of the affine connection
The first three terms on the right-hand side form the Levi–Civita connection,
@@ A.3 The disformation and contortion
A.4 Properties of the torsion, non-metricity, contonsion and the disformation tensors
  • Non-metricity tensor is symmetric in its last two indices:
@@ A.4 Properties of the torsion, non-metricity, contonsion and the disformation tensors
  • Disformation tensor is symmetric in its last two indices:
@@
  • Contonsion is antisymmetric in its first two indices:
@@
  • Difference of the contonsion tensors:
@@
PROOF: The left hand side reads
@@
  • One more difference of the contonsion tensors can be written as:
@@
PROOF: The left hand side reads
@@
PROOF: By definition, the left hand side reads
@@ B.1 The Riemann tensor
Throughout, we use the decomposition derived in Appendix A,
$\Gamma^{\rho}{}_{\mu\nu}={}^{c}\Gamma^{\rho}{}_{\mu\nu}+K^{\rho}{}_{\mu\nu}+L^{\rho}{}_{\mu\nu}$, (93)
where $^{c}\Gamma^{\rho}{}_{\mu\nu}$ is the Levi–Civita connection [7] of $g_{\mu\nu}$, $K^{\rho}{}_{\mu\nu}$ is the contortion, and $L^{\rho}{}_{\mu\nu}$ is the disformation. Substituting (93) into (92) and expanding, one may organize the result as the Levi–Civita curvature plus the terms involving $K$ and $L$. The Levi–Civita Riemann tensor is
@@
where $^{c}\nabla$ denotes the Levi–Civita covariant derivative. Expanding $N^{\rho}{}_{\mu\nu}=K^{\rho}{}_{\mu\nu}+L^{\rho}{}_{\mu\nu}$ in (96), we obtain an explicit expression in terms of contortion and disformation:
@@
which follows from the definition of contortion. With this, one may equivalently present the decomposition in the form used in the main text:
@@
where $\nabla$ is the covariant derivative built from the full connection $\Gamma^{\rho}{}_{\mu\nu}$. The two representations (97) and (99) are equivalent; the former is often cleaner for algebraic manipulations (it uses $^{c}\nabla$), while the latter is convenient when keeping expressions manifestly covariant with respect to the full connection.
@@
The symmetry under the exchange of the first index pair is subtler. In the most general metric–affine case ($Q_{\alpha\mu\nu}\neq 0$), lowering indices involves the full connection and does not commute with covariant differentiation; consequently $R_{\mu\nu\rho\sigma}$ need not to be antisymmetric in $\mu,\nu$. More explicitly, one generally has
@@
Here we used that the disformation tensor is symmetric in its last two indices, $L^{\rho}{}_{\mu\nu}=L^{\rho}{}_{\nu\mu}$, while the contortion satisfies $K_{\rho\mu\nu}=-K_{\mu\rho\nu}$.
@@
Metric-compatible limit.
@@
[…]
$R=^{c}R+2\,\nabla_{\sigma}\!\left(K^{\sigma\mu}{}_{\mu}\right)+K^{\lambda\mu\sigma}K_{\sigma\mu\lambda}+K^{\mu\lambda}{}_{\mu}K^{\sigma}{}_{\lambda\sigma}$. (110)
@@
In this appendix,, we derive the commutator (Ricci) identity for a generic affine connection with torsion and (possibly) nonmetricity. The nonmetricity affects how the indices are raised and lowered, but the commutator formula itself is determined by the torsion and curvature.
@@ C.1 Warm-up: action on a scalar
where the torsion tensor is $T^{\lambda}{}_{\mu\nu}:=\Gamma^{\lambda}{}_{\mu\nu}-\Gamma^{\lambda}{}_{\nu\mu}$. Thus, unlike the Levi–Civita case, the commutator does not vanish even on scalars.
@@ C.2 Action on a covector and a vector
The sign difference between (112) and (113) is the usual one: curvature acts with opposite sign on the covariant versus contravariant indices.
@@ C.3 General Ricci identity
Here each contravariant index contributes a curvature term with a minus sign, while each covariant index contributes a curvature term with a plus sign; the ellipsis indicates the obvious continuation to all remaining indices of the tensor.
@@
In this appendix we derive the generalized Bianchi identities for a generic affine connection with torsion. The derivation is based on the Jacobi identity satisfied by covariant derivatives.
@@ D.1 Jacobi identity
This identity will lead simultaneously to the first and the second Bianchi identities.
@@ D.4 Compact antisymmetric form
In the torsion-free limit $T^{\lambda}{}_{\mu\nu}=0$, these identities reduce to the familiar Bianchi identities of the Riemannian geometry.
@@
Using the splitting of the generic connection
$\Gamma^{\mu}{}_{\nu\sigma}=\ ^{c}\Gamma^{\mu}{}_{\nu\sigma}+K^{\mu}{}_{\nu\sigma}+L^{\mu}{}_{\nu\sigma}$, (125)
and introducing the distortion tensor $D^{\mu}{}_{\nu\sigma}$
@@ E.1 Metric teleparallel connection
In this case the contributions to the distortion tensor are coming from the Levi Civita connection and the torsion tensor. Particularly, one has the equality of the distortion and the contorsion. Namely
@@
Therefore, the Riemann tensor of the Levi-Civita connection in terms of the contorsion tensor reads
@@
Via contraction, one obtains the Ricci tensor as
@@
where the last two term automatically vanishes because of the symmetries. Then
@@
Also one has the following contraction
@@
where the last two term vanish because of the symmetries. Then we have
@@
Renaming the indices it becomes
@@
Finally one ends up with the following expression
@@
The last equation implies that, the Einstein-Hilbert action
@@ E.2 Symmetric teleparallel connection
In this case the distortion is equal to the L-tensor, $D^{\mu}{}_{\nu\sigma}=L^{\mu}{}_{\nu\sigma}$. Therefore, the Riemann tensor of the Christoffel connection can be expressed as
@@
Note that we should be careful in raising the indices because of the non-vanishing of the nonmetricity tensor. Connection is not metric compatible, which means
@@
Remember that $L_{\mu\nu\sigma}$ is symmetric in its last two indices. Hence we can write
@@
Using this result, one can express the scalar curvature of the Levi-Civita connection in terms of $L$, and so the Einstein-Hilbert action. That formulation gives an alternative approach of the general relativity.
@@
This appendix provides the minimal tetrad and the spin–connection [9] framework needed to interpret the teleparallel gravity [10] and to connect the metric–affine wave equations derived in the main text with the gauge–theoretic formulations of the spacetime geometry. No new dynamics are introduced here; the purpose is conceptual unification.
@@ F.3 Torsion and the curvature in differential form language
The torsion and curvature two–forms are defined as
@@ F.4 Teleparallel geometry
The vanishing curvature implies that the spin connection is pure gauge,
@@ F.5 Bianchi identities
In differential–form notation the Bianchi identities take the compact form
@@
These relations are the exterior–calculus counterparts of the component Bianchi identities derived in Appendix D.
This appendix clarifies how the same spacetime geometry may be described equivalently in terms of curvature (Riemannian geometry), torsion (teleparallel geometry), or mixed curvature–torsion structures (metric–affine geometry). This equivalence underlies the wave equations derived in the main text supports the central theme of this work: geometry itself propagates.
@@
This appendix provides a historical and conceptual overview of the geometric ideas underlying the main body of the paper. Our aim is not merely historical completeness, but to clarify why the sequence curvature $\rightarrow$ Ricci identities $\rightarrow$ torsion and nonmetricity $\rightarrow$ tetrads and spin connections $\rightarrow$ teleparallel formulations is both natural and unavoidable.
@@ H.1 Riemann, curvature, and the non-commutativity of derivatives
The modern concept of curvature originates in the seminal work of Bernhard Riemann (1854) [6], who introduced the idea that the geometry is encoded in how distances change infinitesimally. Riemann already understood that the curvature measures the obstruction to Euclidean behavior, but it was only later that this idea was given its algebraic formulation.
@@ H.3 Einstein and the dynamical role of the Bianchi identities
In 1915, Albert Einstein elevated the Riemannian geometry from a mathematical structure to a physical theory of gravitation [5]. By choosing the Levi-Civita connection [7], Einstein ensured that the contracted second Bianchi identity implies
@@ H.6 Teleparallel gravity and Einstein’s alternative viewpoint
Modern Teleparallel Equivalent of general relativity (TEGR) is characterized by
@@ H.7 Symmetric teleparallel gravity and nonmetricity
  • Curvature, torsion, and nonmetricity represent different ways in which covariant derivatives fail to commute or preserve structure. The appendices of this paper make explicit how these descriptions are related and how general relativity emerges as a special, but not unique, geometric choice. For recent studies in various geometries, see [13, 14, 15, 16, 17] and the references therein.
@@ Acknowledgements.
E. Altas is supported by the TUBITAK Grant No. 123F353.

### v2

@@ A.2 Extraction of the affine connection
The first three terms on the right-hand side form the Levi-Civita connection,
@@ A.3 The disformation and contorsion
A.4 Properties of the torsion, non-metricity, contorsion, and the disformation tensors
  • The Non-metricity tensor is symmetric in its last two indices:
@@ A.4 Properties of the torsion, non-metricity, contorsion, and the disformation tensors
  • The disformation tensor is symmetric in its last two indices:
@@
  • contorsion is antisymmetric in its first two indices:
@@
  • Difference of the contorsion tensors:
@@
PROOF: The left-hand side reads
@@
  • One more difference of the contorsion tensors can be written as:
@@
PROOF: The left-hand side reads
@@
PROOF: By definition, the left-hand side reads
@@ B.1 The Riemann tensor
Throughout, we use the decomposition (77) derived in Appendix A. Substituting (77) into (92) and expanding, one may organize the result as the Levi-Civita curvature, along with the terms involving $K$ and $L$. The Levi-Civita Riemann tensor is
@@
where $^{c}\nabla$ denotes the Levi–Civita covariant derivative. Expanding $N^{\rho}{}_{\mu\nu}=K^{\rho}{}_{\mu\nu}+L^{\rho}{}_{\mu\nu}$ in (95), we obtain an explicit expression in terms of contorsion and disformation:
@@
which follows from the definition of contorsion. With this, one may equivalently present the decomposition in the form used in the main text:
@@
where $\nabla$ is the covariant derivative built from the full connection $\Gamma^{\rho}{}_{\mu\nu}$. The two representations (96) and (98) are equivalent; the former is often cleaner for algebraic manipulations (it uses $^{c}\nabla$), while the latter is convenient for keeping expressions manifestly covariant with respect to the full connection.
@@
The symmetry under the exchange of the first index pair is subtler. In the most general metric–affine case ($Q_{\alpha\mu\nu}\neq 0$), lowering indices involves the full connection and does not commute with covariant differentiation; consequently, $R_{\mu\nu\rho\sigma}$ need not be antisymmetric in $\mu,\nu$. More explicitly, one generally has
@@
Here we used that the disformation tensor is symmetric in its last two indices, $L^{\rho}{}_{\mu\nu}=L^{\rho}{}_{\nu\mu}$, while the contorsion satisfies $K_{\rho\mu\nu}=-K_{\mu\rho\nu}$.
@@
Metric-compatible limit
@@
Using $Q_{\alpha\mu\nu}=0$ one may move the metric through the covariant derivative, so $g^{\mu\nu}\nabla_{\sigma}K^{\sigma}{}_{\mu\nu}=\nabla_{\sigma}K^{\sigma\mu}{}_{\mu}$, and therefore
$R=^{c}R+2\,\nabla_{\sigma}\!K^{\sigma\mu}{}_{\mu}+K^{\lambda\mu\sigma}K_{\sigma\mu\lambda}+K^{\mu\lambda}{}_{\mu}K^{\sigma}{}_{\lambda\sigma}$. (109)
@@
In this appendix, we derive the commutator (Ricci) identity for a generic affine connection with torsion and possibly nonmetricity. The nonmetricity affects how the indices are raised and lowered, but the commutator formula itself is determined by the torsion and curvature.
@@ C.1 Warm-up: action on a scalar
where the torsion tensor is $T^{\lambda}{}_{\mu\nu}:=\Gamma^{\lambda}{}_{\mu\nu}-\Gamma^{\lambda}{}_{\nu\mu}$. Thus, unlike the Levi-Civita case, the commutator does not vanish even for scalars.
@@ C.2 Action on a covector and a vector
The sign difference between (111) and (112) is the usual one: curvature acts with the opposite sign on the covariant versus contravariant indices.
@@ C.3 General Ricci identity
Here, each contravariant index contributes a curvature term with a minus sign, while each covariant index contributes a curvature term with a plus sign; the ellipsis indicates the obvious continuation to all remaining indices of the tensor.
@@
In this appendix, we derive the generalized Bianchi identities for a generic affine connection with torsion. The derivation is based on the Jacobi identity satisfied by covariant derivatives.
@@ D.1 Jacobi identity
This identity will lead simultaneously to the first and second Bianchi identities.
@@ D.4 Compact antisymmetric form
In the torsion-free limit $T^{\lambda}{}_{\mu\nu}=0$, these identities reduce to the familiar Bianchi identities of Riemannian geometry.
@@
Using the splitting of the generic connection (77) and introducing the distortion tensor $D^{\mu}{}_{\nu\sigma}$
@@ E.1 Metric teleparallel connection
In this case, the contributions to the distortion tensor come from the Levi Civita connection and the torsion tensor. Particularly, one has the equality of the distortion and the contorsion. Namely
@@
Therefore, the Riemann tensor of the Levi-Civita connection, in terms of the contorsion tensor, reads
@@
By way of contraction, one obtains the Ricci tensor as
@@
where the last two terms automatically vanish because of the symmetries. Then
@@
Also, one has the following contraction:
@@
where the last two terms vanish because of the symmetries. Then we have
@@
By renaming the indices, it has become
@@
Finally, one ends up with the following expression
@@
The last equation implies that the Einstein-Hilbert action
@@ E.2 Symmetric teleparallel connection
In this case, the distortion is equal to the L-tensor, $D^{\mu}{}_{\nu\sigma}=L^{\mu}{}_{\nu\sigma}$. Therefore, the Riemann tensor of the Christoffel connection can be expressed as
@@
Note that we should be careful in raising the indices because of the non-vanishing of the nonmetricity tensor. The connection is not metric compatible, which means
@@
Remember that $L_{\mu\nu\sigma}$ is symmetric in its last two indices. Hence, we can write
@@
Using this result, one can express the scalar curvature of the Levi-Civita connection in terms of $L$, and thus the Einstein-Hilbert action. That formulation provides an alternative approach to general relativity.
@@
This appendix provides the minimal tetrad and the spin–connection [9] framework needed to interpret teleparallel gravity [10] and to connect the metric–affine wave equations derived in the main text with the gauge-theoretic formulations of the spacetime geometry. No new dynamics are introduced here; the purpose is conceptual unification.
@@ F.3 Torsion and the curvature in differential form language
The torsion and curvature two-forms are defined as
@@ F.4 Teleparallel geometry
The vanishing curvature implies that the spin connection is a pure gauge,
@@ F.5 Bianchi identities
In differential-form notation, the Bianchi identities take the compact form
@@
These relations are the exterior calculus counterparts of the component Bianchi identities derived in Appendix D.
This appendix clarifies how the same spacetime geometry may be described equivalently in terms of curvature (Riemannian geometry), torsion (teleparallel geometry), or mixed curvature–torsion structures (metric–affine geometry). The equivalence underlying the wave equations derived in the main text supports the central theme of this work: geometry itself propagates.
@@
This appendix provides a historical and conceptual overview of the geometric ideas underlying the main body of the paper. Our aim is not merely historical completeness, but to clarify why the sequence of curvature $\rightarrow$ Ricci identities $\rightarrow$ torsion and nonmetricity $\rightarrow$ tetrads and spin connections $\rightarrow$ teleparallel formulations is both natural and unavoidable.
@@ H.1 Riemann, curvature, and the non-commutativity of derivatives
The modern concept of curvature originates in the seminal work of Bernhard Riemann (1854) [6], who introduced the idea that geometry is encoded in how distances change infinitesimally. Riemann already understood that curvature measures the obstruction to Euclidean behavior, but it was only later that this idea was given its algebraic formulation.
@@ H.3 Einstein and the dynamical role of the Bianchi identities
In 1915, Albert Einstein elevated Riemannian geometry from a mathematical structure to a physical theory of gravitation [5]. By choosing the Levi-Civita connection [7], Einstein ensured that the contracted second Bianchi identity implies
@@ H.6 Teleparallel gravity and Einstein’s alternative viewpoint
The modern teleparallel equivalent of general relativity (TEGR) is characterized by
@@ H.7 Symmetric teleparallel gravity and nonmetricity
  • Curvature, torsion, and nonmetricity represent different ways in which covariant derivatives fail to commute or preserve structure. The appendices of this paper make explicit how these descriptions are related and how general relativity emerges as a special, but not unique, geometric choice. For recent studies in various geometries, see [22, 23, 24, 25, 26, 27, 28] and the references therein.
@@ Acknowledgements.
We thank Muzaffer Adak for useful discussions and for a careful reading of the manuscript. We are also grateful to Friedrich W. Hehl for an important remark that led to a revision of the table. Professor Metin Gürses has been a constant source of inspiration through his never-fading enthusiasm for physics, his exceptional work ethic, and his generosity as a collaborator and colleague. Having had the privilege of working with him for more than a decade on research and many joint papers, we deeply appreciate both his scientific insight and his kindness as a person. On the occasion of his becoming an emeritus professor at the Bilkent University Department of Mathematics, we celebrate his remarkable contributions to gravitational and mathematical physics and wish him many more years of joyful and fruitful exploration of gravity. E. Altas is supported by the TUBITAK Grant No. 123F353.